### v1

# Language Bottleneck Models: A Framework for Interpretable Knowledge Tracing and Beyond

Abstract
Accurately assessing student knowledge is critical for effective education, yet traditional Knowledge Tracing (KT) methods rely on opaque latent embeddings, limiting interpretability. Even LLM-based approaches generate direct predictions or summaries that may hallucinate without any accuracy guarantees. We recast KT as an inverse problem: learning the minimum natural-language summary that makes past answers explainable and future answers predictable. Our Language Bottleneck Model (LBM) consists of an encoder LLM that writes an interpretable knowledge summary and a frozen decoder LLM that must reconstruct and predict student responses using only that summary text. By constraining all predictive information to pass through a short natural-language bottleneck, LBMs ensure that the summary contains accurate information while remaining human-interpretable. Experiments on synthetic arithmetic benchmarks and the large-scale Eedi dataset show that LBMs rival the accuracy of state-of-the-art KT and direct LLM methods while requiring orders-of-magnitude fewer student trajectories. We demonstrate that training the encoder with group-relative policy optimization, using downstream decoding accuracy as a reward signal, effectively improves summary quality.
1 Introduction
The Challenge of Estimating a Student’s Knowledge State.
A fundamental objective in education is accurately assessing what a student knows, identifying misconceptions, and understanding how their knowledge evolves over time posner1982accommodation ; larkin2012misconceptions ; chen2020impact . Teachers intuitively achieve this through diagnostic reasoning: by observing students’ answers, they infer not merely correctness, but deeper patterns reflecting conceptual mastery or specific misunderstandings. For instance, a student consistently misapplying algebraic properties signals conceptual gaps requiring targeted intervention. Replicating this nuanced interpretability within AI would profoundly enhance personalized education at scale, augmenting human teaching by providing transparent, actionable insights into learners’ knowledge states (khosravi2022explainable, ).
Building algorithms that can learn the knowledge state of students as they interact with educational material is an active field of study known as Knowledge Tracing (KT) (corbett_knowledge_1994, ; shen2024survey, ) (see Appendix A for a detailed introduction to KT). However, accurately modeling a student’s knowledge state in a way that is both predictive and interpretable remains a significant open challenge. Traditional KT methods (piech_deep_2015, ; zhang2017dynamic, ; pandey2019self, ; ghosh2020context, ) rely primarily on statistical, embedding-based representations. Such models lack transparency, representing knowledge states as high-dimensional latent vectors that provide limited interpretability or actionable insights for teachers and learners (Figure 1B). A truly impactful AI assistant in education must go beyond opaque predictions, explicitly articulating what the student understands and what they do not, in an accurate and interpretable manner.
Limitations of Existing LLM-based approaches for KT.
Recent approaches leveraging large language models (LLMs) have shown that LLMs can produce sensible predictions of students’ future behavior when provided with relevant information (li2024explainable, ; kim2024beyond, ), help mitigate the KT’s cold-start problem (lee2024language, ) and be finetuned to improve accuracy on KT tasks (wang2025llm, ). Nonetheless, these LLM-based methods still lack rigorous interpretability, as they either treat the model as a black-box or rely on free-form explanations that are susceptible to hallucination (bender2021dangers, ). Without ensuring faithful and grounded summaries, LLM-generated explanations fall short as reliable educational tools, emphasizing the need for a structured and principled foundation for generating interpretable, reliable, and adaptive student knowledge representations.
Figure 1: Language Bottleneck Models for Knowledge Tracing. The observed questions and answer $\mathcal{D}=\{\mathcal{X},\mathcal{Y}\}$ are caused by a certain knowledge state $\mathcal{S}$ held by the student when answering questions. The language bottleneck aims to approximate this knowledge state in natural language ($\tilde{\mathcal{S}}$). Then, the predicted knowledge state is used to predict both the observed historical data $\mathcal{X}$ and future questions $\mathcal{Y}$.
Knowledge Tracing as an Inverse Problem: Towards Concise and Interpretable Summaries.
Existing Knowledge Tracing methods typically approach student modeling as a forward prediction task, using latent statistical embeddings to predict future responses. However, this approach provides limited insight into the underlying student knowledge. We reframe Knowledge Tracing as an inverse problem: viewing student responses as outcomes of an underlying knowledge state $\mathcal{S}$, manifested through the student’s cognitive process (Figure 1A). The goal is to recover an interpretable representation $\tilde{\mathcal{S}}$ of the student knowledge state. Crucially, answers to downstream questions are expected to arise from the same latent knowledge state—an assumption we discuss further in Section 2.2. A useful representation must therefore reconstruct observed answers, enable prediction of future responses, and remain interpretable and steerable for teachers.
Introducing Language Bottleneck Models (LBMs).
Motivated by the inverse-problem framing described above, we introduce Language Bottleneck Models (LBMs). As illustrated in Figure 1C, LBMs uses two Large Language Models (LLMs) as student knowledge state extractor (encoder) and interpreter (decoder), generating succinct textual summaries meant to capture the student knowledge state. Specifically, LBMs map sequences of student interactions $\mathcal{X}$ to a concise, natural-language summary $\tilde{\mathcal{S}}$. This summary then serves as the sole basis for reconstructing the observed responses $\mathcal{X}$ and accurately predicting future responses $\mathcal{Y}$. By design, this ensures that the generated explanations encompass precisely the information necessary to interpret past behavior and reliably forecast future student performance.
Contributions and Paper Structure.
We make three key contributions:
  • We reframe knowledge tracing as an inverse problem and formalize a minimal-description objective that yields accurate and interpretable summaries of student knowledge states.
  • We introduce Language Bottleneck Models (LBMs)—an encoder-decoder architecture that forces all predictive information through a concise natural-language summary.
  • We empirically validate LBMs on synthetic and real-world datasets against strong KT and direct-prompting baselines, and show how reinforcement finetuning can train the encoder using downstream prediction accuracy.
The remainder of this paper is organized as follows: Section 2 introduces the knowledge tracing problem and reframes it as an inverse problem. Section 3 presents our Language Bottleneck Model (LBM) framework, along with results motivating this approach, and training methods. Section 4 discusses related literature on knowledge tracing and interpretability. Section 5 presents experimental results validating the effectiveness of LBMs. Finally, Section 6 discusses the applicability of LBMs beyond knowledge tracing and concludes with directions for future research.
2 Knowledge Tracing as an Inverse Problem
In this section, we formally introduce the knowledge tracing problem. We then reframe it as an inverse problem—inferring interpretable knowledge states from observed responses, outlining key desiderata for such representations and the conditions under which this formulation is appropriate.
2.1 Knowledge Tracing: Formal Problem Definition
Knowledge Tracing (KT) is fundamentally a sequence modeling problem where the objective is to infer a student’s latent knowledge state based on observed interactions. Formally, let $\mathcal{X}=\{(q_{1},a_{1}),(q_{2},a_{2}),\ldots,(q_{t},a_{t})\}$ denote a sequence of question-answer pairs, where $q_{i}\in\mathcal{Q}$ represents a question from the question space $\mathcal{Q}$, and $a_{i}\in\mathcal{A}$ represents the student’s answer from the answer space $\mathcal{A}$. The goal of KT is to learn a function that predicts the probability of correctly answering a future question $q_{t+1}$ given the historical interaction sequence: $P(a_{t+1}|q_{t+1},\mathcal{X})$.
While this formulation captures the predictive aspect of KT, it fails to address a crucial dimension: the interpretability. An ideal KT system should not only predict accurately but also provide educators with an explicit, human-understandable representation of the student’s knowledge state—such as mastered concepts, knowledge gaps and misconceptions. The importance of interpretability for KT is further discussed in the Appendix.
2.2 Knowledge Tracing as an Inverse Problem: From Black–Box Predictions to Interpretable Summaries
The conventional KT formulation (§2.1) focuses on predictive accuracy but offers limited insight into the underlying knowledge that drives student behavior. We instead frame KT as an inverse problem: in many practical scenarios—such as a quiz completed in a single sitting—a student’s responses can largely be explained by a fixed underlying knowledge state, expressed through the student’s cognitive process of answering questions. Most observed behavior reflects this latent state, with occasional deviations due to noise—such as slips (careless mistakes) or guesses (lucky correct answers), as formalized in Bayesian KT corbett_knowledge_1994 . The goal, then, is to recover an interpretable representation $\tilde{S}$ of the knowledge state that accounts for the observed interactions and supports accurate future predictions.
Framing KT in this way highlights three desiderata. (1) Self-consistency: the inferred summary $\tilde{S}$ should reconstruct the past by making the observed answers in $\mathbf{X}$ likely under a decoder. (2) Predictive sufficiency: $\tilde{S}$ must retain all information necessary to accurately predict future responses to unseen questions. (3) Human interpretability and steerability: the summary should be readable and editable by domain experts, enabling the diagnosis of student misconceptions and manual refinement or augmentation of $\tilde{S}$ to guide future predictions.
When is the inverse formulation appropriate? Although real learning is dynamic, many high–stakes diagnostic situations (end–of–unit quizzes, placement tests, or short tutoring sessions) span minutes rather than weeks. In such windows, it is reasonable to assume a quasi–static knowledge state. In this work, we explicitly focus on scenarios characterized by such quasi–static knowledge states. For longer horizons, the knowledge state might be better modeled as piecewise-static or smoothly evolving $S(t)$. We leave the extension of our framework to dynamically evolving knowledge states for future work.
2.3 Natural Language as the Interface
While a student’s knowledge state is inherently latent, natural language offers a flexible and interpretable medium for expressing it. Unlike fixed concept vocabularies, natural language accommodates open-ended reasoning patterns, misconceptions, and levels of mastery. In this work, we propose to use concise natural language summaries as proxies for student knowledge, designed to capture the latent cognitive state in a way that is both human-readable and predictive. The next section introduces a model family that operationalizes this idea.
3 Language Bottleneck Models
Figure 2: Zero-shot knowledge state encoding compared to ground truth. LLM models with different capabilities are prompted to write a summary of the knowledge state (top left panel) of a student, given 100 observed questions and answers provided in text. The ground truth knowledge state (top right panel) describing the student behavior has three main components: constructs mastered, constructs not mastered, and misconceptions. All three models capture the construct mastery correctly, but are not able to capture any misconception, beside o1 which notices the negative numbers misconception (bottom row). Bottom right symbols correspond to components in Figure 1.
3.1 Formal Definition
To tackle the inverse problem formulation of KT introduced in Section 2, we propose Language Bottleneck Models (LBMs): a two-stage architecture for inferring interpretable and accurate knowledge states from a student’s past interactions.
Let $X_{s}\subseteq X$ denote a subset of observed student responses, and let $Y_{s}\subseteq Y$ denote a set of future questions whose answers the model aims to predict. An LBM is defined by:
  • An encoder $f_{\theta}:X_{s}\rightarrow S$, which maps a history of interactions to a textual summary $S$ of the student’s implicit knowledge state.
  • A decoder $g:S\rightarrow\hat{Y}_{s}$, which uses only the textual summary to predict the student’s answers to new questions.
The explicit use of natural language as a bottleneck encourages the model to produce summaries that balance brevity with completeness. If the language summary omits essential details, predictions of future responses will suffer; conversely, if the summary is overly detailed or redundant, its utility for educators decreases. Thus, this design ensures the resulting textual summary is simultaneously concise, informative, and directly interpretable by humans.
The following sections motivate this approach and outline how LBMs can be trained effectively.
3.2 Motivating Observations
Before detailing our training procedure, we highlight two motivating observations that justify the LBM framework and its constraints.
Observation 1: Summarizing a knowledge state is challenging. Figure 2 compares summaries produced by different LLMs when prompted to describe a student’s knowledge state from 100 question-answer pairs from a synthetic dataset (see Section 5 for details). While all models capture high-level skill mastery, only one correctly identifies a misconception (errors with negative numbers) out of the four existing ones, illustrating that standard summarization approaches often miss crucial latent patterns.
Figure 3: Accuracy on synthetic dataset given ground truth knowledge state summaries.
Observation 2: Given a good summary, capable LLMs can decode with high fidelity. Figure 3 evaluates the performance of different decoder models when given access to the ground truth summary of the student’s latent knowledge state (an example of a ground-truth summary is shown in Figure 2). Stronger models like GPT-4o achieve nearly perfect accuracy ($98\%$), indicating that the bottleneck representation is indeed sufficient to drive effective downstream prediction—provided it encodes the right information.
Together, these observations motivate the inverse formulation: while generating summaries from raw interactions is hard, once obtained, these summaries are powerful inputs for reasoning and prediction. This validates our approach of learning encoders that optimize predictive utility via interpretable, compressed natural language representations.
3.3 Training Objective and Optimization
We train LBMs to produce informative and interpretable summaries of student knowledge states. As demonstrated above, strong decoders such as GPT-4o can achieve near-perfect accuracy when provided with accurate knowledge summaries. We therefore optimize the encoder $f_{\theta}$ to produce such summaries, while freezing the decoder $g$ to a strong LLM during training.
Summary generation and evaluation.
Given a student’s interaction history $X$, we sample a subset $X_{\text{enc}}\subseteq X$ of $N$ interactions to generate a summary $S=f_{\theta}(X_{\text{enc}})$. We evaluate this summary on two disjoint subsets: (i) $X_{s}\subseteq X_{\text{enc}}$: past questions for assessing reconstruction ability; and (ii) $Y_{s}\subseteq Y$: future questions for evaluating predictive accuracy. The summary $S$ is passed to the frozen decoder $g$, which outputs predictions $\hat{X}_{s}=g(S;X_{s})$ and $\hat{Y}_{s}=g(S;Y_{s})$.
Reward function.
The reward for a summary $S$ is defined using a function $\Phi$ that balances multiple criteria:
$\mathcal{R}(S;g)=\Phi\left(\texttt{Acc}(\hat{X}_{s},X_{s}),\;\texttt{Acc}(\hat{Y}_{s},Y_{s}),\;|S|,\;\Omega(S)\right),$ (1)
where $\texttt{Acc}(\cdot)$ measures accuracy, $|S|$ is the summary length, and $\Omega(S)$ enforces structural constraints (e.g., inclusion of a Misconceptions section). The function $\Phi$ balances these components using hyperparameters or indicator functions to enforce constraints as needed.
Optimization via GRPO.
To optimize $f_{\theta}$, we employ Group Relative Policy Optimization (GRPO) shao2024deepseekmath . For each input $X$, the encoder generates $G$ candidate summaries $\{S_{1},\dots,S_{G}\}$, each evaluated using $\mathcal{R}(S_{i};g)$. We compute group-relative advantage and update parameters:
$A(S_{i})=\frac{\mathcal{R}(S_{i};g)-\mu}{\sigma},\quad\nabla_{\theta}\mathbb{E}_{S\sim f_{\theta}(X)}[\mathcal{R}(S;g)]\approx\frac{1}{G}\sum_{i=1}^{G}A(S_{i})\cdot\nabla_{\theta}\log p_{\theta}(S_{i}|X),$ (2)
where $\mu$ and $\sigma$ are the mean and standard deviation of rewards within the group. This approach encourages the encoder to produce summaries that maximize the reward function defined in Eq. 1.
3.4 Steerability of the Estimated Knowledge States
A key advantage of LBMs is the ability for humans to steer the model’s behavior. We identify three mechanisms for human-model interaction (detailed in Appendix):
  • Prompt engineering the encoder. Directly shaping summary generation through input instructions.
  • Steering via reward signals. Incorporating human preferences during encoder training through the $\Omega(S)$ term in Eq. 1.
  • Augmenting with student-specific information. Educators can supplement the model with additional knowledge not present in observed data—either by augmenting encoder inputs or by editing the generated summary before decoding. This enables integration of recent classroom observations or specific misconceptions identified through in-person interactions.
4 Related Work
In this section, we review the literature related to our proposed approach, spanning traditional knowledge tracing, recent advances in LLM-based student modeling, and concept bottleneck models (CBMs), and text summarization models.
Traditional Knowledge Tracing
Traditional knowledge tracing (KT) methods have long been used to model student learning and predict future performance shen2024survey . Early works such as Bayesian Knowledge Tracing (BKT) corbett_knowledge_1994 model student mastery as a probabilistic process, while more recent methods like Deep Knowledge Tracing (DKT) piech_deep_2015 and Attentive Knowledge Tracing (AKT) ghosh2020context uses an attention-based architecture to learn contextual representation of questions and student knowledge states. Despite their strong predictive performance, these models represent a student’s knowledge state as an abstract, high-dimensional latent vector, which poses significant challenges in interpretability and actionable feedback for educators.
LLM-Based Knowledge Tracing
Recent studies have begun integrating Large Language Models (LLMs) into the KT framework. For example, li2024explainable demonstrated that LLMs are able to make sensible predictions about student responses when prompted with adequate information. Other works lee2024language ; kim2024beyond have studied how LLMs can help mitigate the cold-start problem compared to traditional KT approaches, while wang2025llm demonstrated state-of-the-art performances in KT by combining LLMs with sequence interaction models. However, these methods generally remain opaque: they either treat the LLM as a black-box, or rely on model-generated explanations that are susceptible to hallucination (bender2021dangers, ).
Concept Bottleneck Models
Concept Bottleneck Models (CBMs) koh2020concept improve interpretability through human-understandable concept activations as intermediates, with extensions exploring unsupervised concept learning oikarinen2023label , test-time interventions shin2023closer , and theoretical analyses of concept set design luyten2024theoretical . However, CBMs typically rely on finite predefined concept sets, limiting their applicability to complex tasks like knowledge tracing. Recently, yamaguchitoward introduced Explanation Bottleneck Models (XBMs), which use textual rationales as intermediates for vision classification, justifying a single known label per input. While our Language Bottleneck Models (LBMs) adopt this language bottleneck concept, they differ fundamentally. Unlike XBMs’ instance-specific, task-specific rationales, LBM summaries aim to capture an implicit knowledge states—as emphasized by our inverse problem formulation—and generalize to future, unknown questions. These requirements necessitate holistic, adaptable summaries rather than XBMs’ task-specific rationales, leading us to introduce Language Bottleneck Models (LBMs) as a distinct framework with broader applicability.
Contrast with Text Summarization Models
Recent text summarization models such as BART lewis2019bart , T5 raffel2020exploring , and PEGASUS zhang2020pegasus effectively produce concise summaries based on explicitly available textual content. However, these approaches differ fundamentally from knowledge tracing, where summaries must infer implicit student knowledge states not directly observable in the input. Standard summarization metrics (e.g., ROUGE, BLEU) rely on explicit reference summaries, making them unsuitable for evaluating latent knowledge inference tasks. In contrast, our Language Bottleneck Models generate textual summaries of the student’s implicit knowledge state, optimized for predictive accuracy on downstream questions rather than syntactic overlap with the observed input.
5 Experiments
We evaluate the effectiveness of our Language Bottleneck Models (LBM) in the context of Knowledge Tracing (KT), both with a controlled synthetic KT task involving basic arithmetic operations, and on real-world knowledge tracing datasets.
5.1 Experimental Setup
Datasets We evaluate on one synthetic and one real-world source. Synthetic arithmetic simulates learners answering addition, subtraction, multiplication, and division items; each student is assigned mastered skills, unmastered skills, and systematic misconceptions. The Eedi dataset wang2020diagnostic contains millions of question–answer pairs from UK secondary-school mathematics. More details about these datasets are provided in the Appendix.
Models The main experiments use GPT-4o and GPT-4o-mini, together with open-source Qwen 2.5 (3B, 7B) and Gemma 3 (12B, 30B) checkpoints. Unless noted, we run the instruct variants of each open-source model; hyper-parameters and prompt templates are provided in the Appendix.
Baselines We compare LBMs against (i) four standard knowledge-tracing methods—DKT piech_deep_2015 , DKVMN zhang2017dynamic , SAKT pandey2019self , and AKT ghosh2020context —implemented with the pyKT library liupykt2022 ; and (ii) direct prompting, where each LLM predicts answers from the full student history without a bottleneck. Training details for all baselines are given in Appendix.
5.2 Results
5.2.1 LBM vs Direct Prompting
Table 1: Accuracy results for Direct and LBM methods. Bold indicates LBM accuracy is no more than 2% below the Direct baseline.
synthetic eedi Direct LBM $\Delta$ Direct LBM $\Delta$ Qwen2.5-3B 0.525 0.627 0.102 0.347 0.372 0.025 Qwen2.5-7B 0.655 0.646 -0.009 0.647 0.583 -0.065 gemma-3-12b 0.790 0.791 0.001 0.580 0.625 0.045 gemma-3-27b 0.785 0.794 0.009 0.670 0.650 -0.020 gpt-4o-mini 0.795 0.782 -0.013 0.655 0.645 -0.010 gpt-4o 0.840 0.825 -0.015 0.703 0.682 -0.020
Figure 4: Evolution of accuracy with bottleneck length.
Table 1 compares the accuracy of LBM models to direct prompting variants across different LLMs. Despite the information bottleneck enforced by LBMs, they remain within $2\%$ of the accuracy obtained via direct prompting. Figure 4 demonstrates that this gap decreases for longer bottlenecks, highlighting a trade-off between conciseness and predictive accuracy.
Encoder/decoder variants
We evaluate LBMs with different models as encoders and decoders (results shown in the Appendix). Using the strongest model (gpt-4o) as encoder with other models from Table 1 as decoders yields accuracies $5-10\%$ higher than when the stronger model is used as the decoder. This confirms our hypothesis that extracting relevant information for the summary (encoding) is more challenging than predicting future answers given a summary (decoding).
5.2.2 LBM vs Knowledge Tracing Methods
Figure 5 compares the performance of LBM models to traditional Knowledge Tracing methods on the synthetic dataset. LBMs with strong LLMs achieve comparable accuracy to KT methods with significantly less training data. Since KT methods rely on statistical patterns, they require substantially more observations before reaching strong predictive power, while LBMs demonstrate effective generalization with minimal examples.
Figure 5: Accuracy of LBM vs Knowledge Tracing models on the Synthetic dataset. The x axis represents the total number of question seen by the model, ie the product of number of questions observed per student $T$ and the number of students available for training $N$ (note that N=1 for LBM methods). The y axis is the average accuracy 200 test student and 4 new questions to predict per student. The pre-trained models indicated for LBM are used as both Encoder and Decoder, with chain-of-thought prompting. Error bars report the standard error calculated across 10 runs for N=10 students and 3 runs otherwise.
5.2.3 Training LBM Encoders
As discussed in Section X, downstream accuracy serves as a reward signal to train encoders to produce increasingly accurate summaries. We demonstrate this by training a Gemma3-12B encoder with GRPO alongside a frozen Gemma3-27B decoder. Figure 6 shows the encoder progressively improving summary quality, quickly outperforming gpt-4o. The figure illustrates this progression through an example student who mastered all constructs except subtraction and fails any multiplication involving 9. The initial summary contains inaccuracies and misses this systematic misconception, while the final summary successfully captures the student’s complete knowledge state.
Figure 6: Training the encoder with the decoding accuracy on the synthetic dataset. Evolution of the test accuracy as a Gemma3-12B encoder is trained as described in Section 3.3, with a fixed Gemma3-27B decoder. Trained on $800$ training students and tested on $200$ students, with $|\mathcal{X}|=50$ questions per input trajectory and $|\mathcal{Y}|=20$ questions to predict per student. The bottom row show the evolution of the generated summary over the course of training for an example student. Text is colored green (exact), orange (approximate), or red (false) based on ground-truth.
Accuracy gains from training are stronger on harder cases
When stratifying students by difficulty (measured by number of misconceptions), we find training yields larger accuracy improvements for complex cases (see table in the Appendix). The model shows 10-11% increased accuracy relative to gpt-4o for students with 2 or 3+ misconceptions, compared to just 5-7% for students with 0-1 misconceptions. This suggests the encoder becomes particularly better at handling complex knowledge states.
5.2.4 Steering LBM behavior
We demonstrate multiple steering strategies described in Section 3.4 (details in Appendix). We find that providing explicit misconception information during encoder training produces significantly stronger learning effects than adding the same information at the decoder stage. This suggests the encoder leverages this additional context to better interpret patterns in student responses. We also show the encoder can be steered during training to explicitly mention misconceptions through reward signals, and illustrate how supplementing the bottleneck with information not present in the input can significantly improve decoder accuracy.
6 Discussion
Core contributions
Framing Knowledge Tracing as an inverse problem shifts focus from fitting statistical regularities to understanding the underlying reasons for student responses. Language Bottleneck Models operationalize this view by compressing interaction histories into explicit natural-language summaries that maintain high predictive power while being both interpretable and steerable by domain experts.
Why can’t we just prompt GPT-4o directly?
The split encoder-decoder architecture of LBMs offers three key advantages over direct LLM prompting: it creates a global student model with a single latent summary shared across all predictions rather than isolated per-question reasoning; it ensures faithful summaries through a closed-loop decoding objective that penalizes non-predictive summaries; and it provides an explicit interface layer that teachers can read, steer and intervene on.
Wider applicability
LBMs extend beyond education to any task requiring compact, human-readable summaries with predictive power. The minimal ingredients needed are: (1) a sequence of observations about an entity, (2) a need to predict future behaviors of that entity, and (3) value in having interpretable representations. For example, clinical decision support could distill patient data into textual state descriptions that forecast outcomes while remaining auditable; preventive maintenance could compress sensor logs into explanations predicting machine failure; customer success teams could summarize interaction histories to forecast churn.
Limitations
LBMs face several constraints: context length limitations imposed by underlying LLMs, which could be addressed through iterative summarization; the requirement for textual question content rather than just identifiers, limiting applicability to some KT datasets; and the substantial computational cost of running LLMs. We discuss limitations extensively in the Appendix.
Future work
Several promising extensions include: implementing iterative encoding to process longer inputs through sequential chunking; developing active sensing capabilities to select questions that maximize information about students’ knowledge states; adapting LBMs for evolving knowledge states to capture learning and forgetting over time; expanding beyond question-answer inputs to leverage richer data sources; and combining with pedagogical-alignment techniques like LearnLM (team2024learnlm, ). We discuss these points in more details in the Appendix. By bridging the gap between predictive power and human interpretability, Language Bottleneck Models represent a significant step toward AI systems that can genuinely augment human expertise in education and beyond.
Appendix A Introduction to Knowledge Tracing
In this section, we provide a more detailed introduction to Knowledge Tracing. For recent comprehensive reviews of the field, see [1, 25].
Knowledge Tracing (KT) is the task of modeling students’ knowledge states and predicting their performance on future questions based on their historical interactions with educational content. The goal is to dynamically track the acquisition, retention, and application of knowledge as students engage with learning materials.
A.1 Formal Problem Definition
Knowledge Tracing (KT) is fundamentally a sequence modeling problem where the objective is to infer a student’s latent knowledge state based on observed interactions. Formally, let $\mathcal{X}=\{(q_{1},a_{1}),(q_{2},a_{2}),\ldots,(q_{t},a_{t})\}$ denote a sequence of question-answer pairs, where $q_{i}\in\mathcal{Q}$ represents a question from the question space $\mathcal{Q}$, and $a_{i}\in\mathcal{A}$ represents the student’s answer from the answer space $\mathcal{A}$. The goal of KT is to learn a function that predicts the probability of correctly answering a future question $q_{t+1}$ given the historical interaction sequence: $P(a_{t+1}|q_{t+1},\mathcal{X})$.
In practice, each interaction $(q_{i},a_{i})$ in the dataset typically contains richer information. Questions $q_{i}$ often include metadata about the underlying knowledge constructs they assess (e.g., specific skills, concepts, or learning objectives). For example, a mathematics question might be tagged with constructs like "quadratic equations" or "algebraic factoring." Similarly, observed answers $a_{i}$ can vary in granularity—from binary correctness indicators to specific answer choices in multiple-choice questions, or complete textual/code responses for open-ended questions. Many KT datasets also capture additional contextual information such as response time or self-reported confidence scores, which can be leveraged by KT models.
The prediction target can also extend beyond binary correctness. Models may predict specific answer choices a student will select (see for example Task 2 of Eedi’s NeurIPS 2020 Education Challenge [30]), enabling more nuanced understanding of misconceptions. In domains like programming education, models might predict the exact code a student will write [14], capturing finer-grained knowledge states.
A.2 Knowledge Tracing Paradigms
Bayesian Knowledge Tracing (BKT).
Introduced by [5], BKT models student knowledge as a set of binary latent variables corresponding to specific skills or concepts. For each skill, a student is either in a "known" or "unknown" state. BKT employs a Hidden Markov Model to update these probabilities based on observed student responses, accounting for learning, guessing, and slipping in its probabilistic framework. This approach provides interpretable outputs, with each concept having a probability of mastery, but is limited in modeling the complex interdependencies between skills.
Deep Learning Approaches.
Deep Knowledge Tracing (DKT) [20] marked a paradigm shift by using recurrent neural networks to model student knowledge as continuous latent representations. Unlike BKT’s discrete skill probabilities, DKT captures complex patterns and skill relationships through high-dimensional embeddings, significantly improving prediction accuracy. Subsequent models have enhanced this approach, including: Dynamic Key-Value Memory Networks (DKVMN) [34] which incorporate memory components for better concept tracking; Self-Attentive Knowledge Tracing (SAKT) [19] and Attentive Knowledge Tracing (AKT) [6] which leverage attention mechanisms to capture long-range dependencies; and Graph-based Knowledge Tracing models [17, 33] which incorporate structural information into knowledge representations.
A.3 Current Challenges
Interpretability vs. Accuracy Trade-off.
Deep learning approaches typically achieve higher predictive accuracy but at the cost of interpretability. Their latent representations are opaque high-dimensional vectors that provide limited insight into specific concepts a student has mastered or misconceptions they hold. This limits the usability of these models in practice.
Why do we need interpretability?
At first glance, predictive accuracy alone might seem sufficient for KT—an accurate model could, for example, analyze student responses across many questions to identify overall strengths and weaknesses. However, even highly accurate models can yield misleading or fragile predictions if the underlying reasoning is opaque. For instance, [22] showed that predictive accuracy alone might obscure models relying on superficial patterns like repeated questions, limiting practical usability. Thus, without interpretability, accurate models may fail to meaningfully support educators in diagnosing and intervening in students’ learning processes.
Cold-start Problem.
KT systems struggle with new students or questions where historical data is limited or non-existent: for any new question added to the dataset, traditional models require the observation of answers from many students to learn an accurate latent representation for this question. This "cold-start" problem poses significant challenges for the practical deployment of KT systems, particularly in dynamic educational environments where new content and students are regularly introduced.
Appendix B Extended Results
B.1 Encoder/Decoder Variants (Section 5.2.1)
Table B1 shows the accuracy of different combinations of encoder-decoder models on the Synthetic dataset. The top row shows the result of using the strongest model evaluated (gpt-4o) as both encoder and decoder. Then, we vary either the encoder or decoder part of the LBM across other LLMs, and report the resulting accuracy. A clear pattern which emerges is that the resulting LBMs are stronger when gpt-4o is used as the encoder instead of the decoder. This implies that the task of accurately capturing knowledge state information in the bottleneck is harder than predicting answers to future questions when provided with a knowledge state summary.
Table B1: Performance of different LBM encoder-decoder combinations on the Synthetic dataset.
Encoder Decoder Accuracy
Strongest model gpt-4o gpt-4o 0.809
Strongest gpt-4o gpt-4o-mini 0.821
model gpt-4o google/gemma-3-12b-it 0.795
as gpt-4o Qwen/Qwen2.5-7B-Instruct 0.765
encoder gpt-4o Qwen/Qwen2.5-3B-Instruct 0.715
Strongest gpt-4o-mini gpt-4o 0.765
model google/gemma-3-12b-it gpt-4o 0.775
as Qwen/Qwen2.5-7B-Instruct gpt-4o 0.666
decoder Qwen/Qwen2.5-3B-Instruct gpt-4o 0.649
B.2 Traditional KT methods on the Eedi dataset
We run the four benchmark KT methods on the Eedi dataset with the same experimental setup as for Table 1 in the main paper (see Section D.2.2 in this appendix for dataset details). Table B2 shows the resulting accuracies, alongside the accuracy found with Direct and LBM models for different LLMs, reproduced from Table 1. The best KT model (AKT) slightly outperform the best LBM and Direct models with gpt-4o, while other KT models reach similar accuracies to LBM and Direct models with gpt models and Gemma3-27B.
Table B2: Average accuracy on the Eedi dataset for different model types. For each student, the accuracy is computed on the last four questions of the trajectory. KT models show mean $\pm$ SEM across 10 runs. Due to computational constraints, Direct and LBM results are reported from single runs.
Model Accuracy
KT Models
DKT $0.647\pm 0.005$
DKVMN $0.669\pm 0.004$
SAKT $0.680\pm 0.007$
AKT $0.720\pm 0.008$
Direct
Qwen2.5-3B 0.347
Qwen2.5-7B 0.647
Gemma3-12B 0.580
Gemma3-27B 0.670
GPT-4o-mini 0.655
GPT-4o 0.703
LBM
Qwen2.5-3B 0.372
Qwen2.5-7B 0.583
Gemma3-12B 0.625
Gemma3-27B 0.650
GPT-4o-mini 0.645
GPT-4o 0.682
B.3 Training LBM Encoders - Difficulty stratification (Section 5.2.3)
To investigate whether the accuracy gains seen during LBM training (cf Figure 6 in the main paper) vary across students in the dataset, we stratify students according to how many misconception they hold: 0, 1, 2 or 3+. Since each misconception represent additional "irregularities" in the student’s response pattern beyond simple mastery of constructs, this effectively stratifies different difficulty levels across students. Table B3 shows the change in accuracy relative to the gpt-4o baseline across difficulty levels.
Table B3: Relative difference in accuracy between trained Gemma3-12B and the GPT-4o baseline, with students grouped by number of misconceptions.
Accuracy (mean ± std) Relative Diff. (%)
# of Misconceptions Baseline Pre-training Post-training Pre Post $\Delta$
0   ($N=22$) 0.98 ± 0.05 0.92 ± 0.12 0.99 ± 0.05 -5.80 0.90 6.70
1   ($N=17$) 0.89 ± 0.09 0.87 ± 0.15 0.91 ± 0.09 -2.00 3.00 5.00
2   ($N=23$) 0.81 ± 0.16 0.80 ± 0.12 0.89 ± 0.08 -0.80 10.20 11.00
3   ($N=135$) 0.77 ± 0.12 0.74 ± 0.14 0.82 ± 0.11 -4.80 5.70 10.50
B.4 Steering LBM behavior
B.4.1 Steering via reward signals
To demonstrate the possibility to steer LBMs’ behavior via the reward signal, we consider an example use-case where a teacher would like the model to pay particular attention to potential misconceptions held by the student. Following the reward-shaping framework of Section 3.3, we augment the training objective with an additional term that explicitly encourages the model to surface misconceptions. Concretely, we set $\Omega(S)=\mathbbm{1}\!\bigl{[}\text{``misconception''}\in\tilde{\mathcal{S}}\bigr{]}$, where $\tilde{\mathcal{S}}$ is the textual bottleneck emitted by the LBM for student state $S$. This binary reward is added to the accuracy term and optimized with GRPO. Figure B1 confirms that the policy quickly internalizes this incentive: after only a handful of training steps, the proportion of bottlenecks that explicitly mention a misconception goes from less than 80% to >95%, demonstrating that the reward function provides an effective lever for shaping higher-level pedagogical behavior.
Figure B1: Presence of the word "misconception" in the LBM’s bottlenecks during training of the encoder model steered towards mentioning student misconceptions via the reward signal.
B.4.2 Augmenting LBMs with student-specific information to complement input data
To demonstrate how an LBM can be actively steered with additional information we run the following ablation on the Synthetic dataset: 1. Each student trajectory originally probes four constructs. For a given trajectory, we remove every question linked to one construct leaving the input data intentionally incomplete. 2. We then run the input data through the encoder and inject a single teacher-supplied sentence describing the student’s mastery of the missing construct directly into the model’s bottleneck representation (e.g., "The student has mastered addition except in the event of misconceptions.).
We repeat this procedure four times—once for each construct—running the LBM both with and without the additional sentence, and report the results Table B4. Naturally, the models complemented with the additional information of the student’s mastery of the construct missing in the input data outperform models provided only with the incomplete data. This small experiment illustrates a key advantage of LBMs: because the LBM compresses evidence into a text-based bottleneck, it can be complemented with additional information absent from the input data.
Table B4: Accuracy of a gpt-4o-based LBM on the Synthetic dataset while removing one construct from the input data, with and without providing information about the missing construct in the bottleneck. Each run is repeated across all four constructs; mean and standard deviation are reported.
Without additional bottleneck information With additional bottleneck information
0.749 ± 0.007 0.791 ± 0.013
B.4.3 Augmenting with student-specific information to assist LBM training
Finally, we demonstrate how providing additional information to the LBM can assist with the training process. We train two identical LBMs while providing 2 misconceptions either to the encoder as part of the input data, or to the decoder as part of the bottleneck. The models are trained for one epoch on $800$ students. We then evaluate the resulting models without any additional information provided. Table B5 shows the accuracy before/after training for both models. The model where additional information was provided to the encoder reaches $84\%$ accuracy after one training compared to $80\%$ when it is provided to the decoder. This suggest that the additional information facilitates training of the encoder.
Table B5: Relative difference in accuracy between trained Gemma3-12B and the gpt-4o baseline, with students grouped by number of misconceptions.
Before training After 1 epoch training
Information in X 0.802 ± 0.001 0.840 ± 0.004
Information in S 0.794 ± 0.002 0.799 ± 0.006
B.5 Example bottlenecks on the Eedi dataset
Figure B2 shows example bottlenecks produced by a gpt-4o-based LBM for a student in the Eedi dataset. The input data is composed of 30 questions from various constructs, and the decoder predicts 4 test questions. As shown Figure 4 in the main paper, a shorter bottleneck of 128 tokens constrains the expressive power of the model, and in this example the resulting predictions fail on two of the test questions. A longer bottleneck of 256 or 512 tokens allows for more nuance and details in the bottleneck, leading to the decoder correctly predicting all four test questions for this student.
Figure B2: Example bottlenecks produced by an LBM with gpt-4o as both encoder and decoder for a student in the Eedi dataset, with varying bottleneck lengths. The input data provided to the encoder is composed of 30 questions across various constructs, and the decoder predicts 4 test questions using the bottleneck.
Appendix C Extended Discussion
C.1 Extended Discussion
Knowledge Tracing as an Inverse Problem
The reframing of Knowledge Tracing as an inverse problem represents a fundamental shift in how we conceptualize student modeling. Traditional KT methods adopt a forward modeling approach, where the goal is to learn a mapping from observed interactions to future performance predictions. While effective for prediction, this treats the student’s knowledge state as an opaque latent variable that serves primarily as an intermediate computational step. In contrast, our inverse problem formulation explicitly seeks to recover the underlying knowledge state that generates observed behaviors, making interpretability the primary objective rather than an afterthought. This aligns with educational assessment principles, where the main objective is to understand what a student knows, rather than merely predicting future behavior.
Parallel with Kolmogorov Complexity
Our inverse problem formulation of Knowledge Tracing draws a natural parallel with algorithmic information theory. The Kolmogorov complexity $K(x)$ of a string $x$ is defined as the length of the shortest program that outputs $x$ when run on a universal Turing machine. Traditional KT methods focus on learning statistical regularities between questions and responses, without looking for parsimonious explanations for observed behavior. In contrast, LBMs explicitly search for minimal natural language descriptions that can both reconstruct past interactions and predict future responses. The knowledge state summary $\mathcal{S}$ is akin to the minimal program, and the decoder LLM functions is akin to the universal interpreter that unpacks $\mathcal{S}$ to generate the observed question-answer patterns. This connection suggests that effective knowledge state summaries should capture the algorithmic essence of student behavior—the underlying "program" of knowledge and misconceptions that generates observable responses—rather than merely fitting surface-level patterns. While true Kolmogorov complexity is uncomputable, LBMs approximate this ideal through the natural language bottleneck constraint, encouraging summaries that balance compression with predictive power.
Computational Architecture and Design Choices
The encoder-decoder architecture reflects a deliberate separation of concerns: extraction versus interpretation of knowledge states. Our experiments demonstrate that these two tasks have different intrinsic complexity, with encoding (summary generation) being more challenging than decoding (prediction from summaries). This finding has practical implications for model selection and computational resource allocation. The use of GRPO for encoder training represents a novel application of reinforcement learning to interpretable AI, where the reward signal directly measures the downstream utility of explanations rather than their surface-level quality.
C.2 Extended Limitations
Context Length Constraints
A practical limitation of LBMs is the context length restrictions of current LLMs. Eventhough modern models can handle thousands of tokens, comprehensive student histories in real educational settings can easily exceed these limits. As student trajectories grow more complex and include more information, the context required for the encoder might go beyond what current LLMs are capable of. A natural solution could be an iterative encoding process, where the encoder iterates over the text bottleneck while going over the input data by windows.
Computational Cost and Resource Requirements
Since LBMs involve two LLMs as the encoder and decoder, they are typically more computationally intensive to run than traditional KT methods. Each inference call with LBMs requires two LLM calls (encoding and decoding) with extensive context windows. Training LBMs with GRPO is particularly costly, as it requires iteratively generating and evaluating multiple candidate summaries per training example. This cost structure may limit the practical deployment of LBMs to high-stakes educational contexts where the interpretability benefits justify the computational expense.
Dependence on Textual Question Content
Many existing KT datasets provide only question identifiers rather than full question text. This limits the direct applicability of LBMs, as they require the questions content in order to generate meaningful summaries about student knowledge. While question content is increasingly available in modern educational platforms, this dependency creates a barrier to applying LBMs to historical datasets or systems that rely primarily on item response theory frameworks.
Quasi-Knowledge State Assumption
The quasi-static knowledge state assumption underlying our inverse problem formulation, while reasonable for short diagnostic sessions, does not cover longer time horizons. Real learning involves continuous knowledge acquisition, forgetting, and misconception evolution that our current framework cannot capture. This limitation restricts current LBMs to diagnostic rather than contexts of assessment throughout the learning process. The assumption also fails to account for contextual factors (fatigue, motivation, external stressors) that can significantly impact student performance within even short sessions.
C.3 Extended Future Work
Iterative Encoding Strategies
A natural extension to address context length limitations involves iteratively applying the encoder while chunking input data. At each step, the encoder would process a portion of the interaction history $X$ along with one or more previously generated bottleneck summaries, creating an updated summary that incorporates new information from the latest chunk. This approach could maintain both detailed recent context and compressed historical patterns, enabling LBMs to handle arbitrarily long student trajectories through sequential refinement of knowledge state representations.
Active Learning and Question Selection
Building on iterative encoding, active learning strategies could provide encoders with the most informative input questions. Rather than processing all available interactions, the system could strategically select questions that maximize information gain about uncertain aspects of student knowledge. This could leverage uncertainty quantification methods or use the encoder LLM itself to recommend the most diagnostically valuable questions. Such active sensing would improve both efficiency and diagnostic power by focusing computational resources on the most insightful student responses.
Dynamic Knowledge State Modeling
Extending LBMs to handle evolving knowledge states represents a significant next step necessary to account for progressive learning and forgetting effects. A natural relaxation of the static knowledge state assumption involves "piecewise-constant" knowledge states that can evolve between question sessions but remain static within sessions. This extension poses exciting challenges in developing training objectives that balance within-session consistency with between-session learning dynamics, potentially requiring new approaches to temporal modeling in natural language representations.
Multimodal Input Integration
Question-answer pairs represent a relatively narrow information source for inferring student knowledge states. Richer data sources such as student-tutor interactions, self-reported explanations for answers, timing patterns, or hint usage could provide deeper insights into student understanding. While LBMs naturally extend to text-based inputs, future work should investigate optimal strategies for combining these varied data sources and evaluate their relative contributions to knowledge state inference accuracy and interpretability.
Pedagogical Alignment
Recent work like LearnLM [28] demonstrates the potential for making LLMs more pedagogically aligned. Incorporating similar pedagogical principles into LBM components could help encoders better interpret educational interactions and generate summaries that align with expert teaching practices. This might involve training on educator-annotated examples, incorporating educational taxonomies into summary structure, or using reward functions that emphasize pedagogically relevant aspects of student knowledge states. Such alignment could bridge the gap between computational convenience and educational validity.
Appendix D Experimental Details
D.1 Models
LLM models
We evaluate the following closed- and open-source models:
  • GPT-4o and GPT-4o-mini[2];
  • Qwen2.5-3B and Qwen2.5-7B[29];
  • Gemma3-12B and Gemma3-27B[27].
For open-source models we extract the activation logits of the “Yes” and “No” tokens and return the higher value. For closed-source models we prompt for a “Yes” or “No” answer and parse the text output.
Knowledge Tracing models
We evaluate the following four Knolwedge Tracing models:
  • DKT [20]
  • DKVMN [34]
  • AKT [6]
  • SAKT [19]
For each one of these models, we use the pyKT implementation [15] with default hyperparameters.
D.2 Datasets
D.2.1 Synthetic
Our synthetic dataset simulates students answering basic arithmetic problems. Each student is characterised by (i) mastered skills, (ii) unmastered skills, and (iii) systematic misconceptions. Questions are arithmetic operations—addition, subtraction, multiplication, or division (rounded to the nearest integer)—between two operands drawn from $[0,15]$ for addition/subtraction or $[1,10]$ for multiplication/division.
Misconception pool.
For every student we sample misconceptions uniformly at random from:
  • forgets to carry in addition;
  • fails multiplications involving the number $x$ with $x\sim\mathcal{U}(6,9)$;
  • fails any operation involving the number $x$ with $x\sim\mathcal{U}(6,9)$;
  • fails whenever an operand $>10$;
  • always rounds division results down;
  • fails with negative numbers.
Generation parameters.
  • Number of students: $2000$;
  • Number of questions: $5000$;
  • Questions answered per student: $210$.
Figure D1 shows the histogram of the number of misconceptions per student across the $2000$ students of the dataset.
Figure D1: Distribution of the number of misconceptions per student in the Synthetic dataset.
D.2.2 Eedi
The Eedi dataset is analogous to the one publicly shared via the NeurIPS 2020 Education Challenge [30] organised by Eedi, but additionally includes the text of each question.
Although many trajectories in the Eedi dataset span several days or months, we retain only single-session sequences to satisfy the quasi-static-knowledge assumption:
  • minimum response time: $5\,\mathrm{s}$;
  • maximum gap between answers: $3\,\mathrm{min}$;
  • minimum trajectory length: $40$ questions.
This filtering yields a total of $623$ individual trajectories with 40+ questions.
D.3 Results
D.3.1 Table 1
Synthetic dataset
  • $|\mathcal{X}|=50$
  • $|\mathcal{Y}|=4$
  • Test students: $200$
  • Bottleneck size: $128$ tokens
  • Chain-of-thought prompting: No
Eedi dataset
  • $|\mathcal{X}|=30$
  • $|\mathcal{Y}|=4$
  • Test students: $100$
  • Bottleneck size: $512$ tokens
  • Chain-of-thought prompting: No
D.3.2 Figure 4
Same set-up as Table 1; only the bottleneck size varies (see $x$-axis).
D.3.3 Figure 5
Latent Bottleneck Models (LBMs)
  • $|\mathcal{X}|=50$, $|\mathcal{Y}|=4$
  • Test students: $200$
  • Bottleneck size: $256$ tokens
  • Chain-of-thought prompting: Yes
  • Encoder & decoder: identical (varies by row in the table)
Knowledge-tracing baselines DKT [20], DKVMN [34], AKT [6], and SAKT [19] are evaluated using the pyKT implementation [15], modified to compute accuracy on $|\mathcal{Y}|=4$ questions, in order to be comparable with our LBM models. For each method we keep the pyKT default hyperparameters.
D.3.4 Figure 6
  • Dataset: Synthetic
  • $|\mathcal{X}|=50$, $|\mathcal{Y}|=20$
  • Train / test students: $800/200$
  • Bottleneck size: $128$ tokens
  • Chain-of-thought prompting: No
  • Encoder (trained): Gemma3-12B
  • Decoder (frozen): Gemma3-27B
Training hyper-parameters.
[…]
  • $G=5$
  • Learning $\text{Learning rate}=1\times 10^{-4}$
  • $\beta=0.04$
  • Optimiser: Adam (adamw_torch, default settings)
  • LoRA configuration
    • $r=16$, $\alpha=16$
    • target_modules = ["q_proj", "k_proj", "v_proj", "o_proj", "gate_proj", "up_proj", "down_proj"]
    • Dropout: $0.05$
D.3.5 Encoder–Decoder Variants
Identical to the Table 1 set-up, except for the choice of encoder and decoder LLMs.
D.3.6 Difficulty Variations
Synthetic students are stratified by the number of misconceptions held: 00, $1$, $2$, or $3+$.
D.3.7 Steering Experiments
Same dataset and model parameters as for Figure 6.
Training hyper-parameters.
[…]
  • $G=4$
  • Learning $\text{Learning rate}=1\times 10^{-4}$
  • $\beta=0.04$
  • Optimiser: Adam (adamw_torch, default settings)
  • LoRA configuration
    • $r=16$, $\alpha=16$
    • target_modules = ["q_proj", "k_proj", "v_proj", "o_proj", "gate_proj", "up_proj", "down_proj"]
    • Dropout: $0.05$
D.4 Prompts
Below are the different prompts used to query LLMs in our experiments.
1template: |
2 Please produce a concise summary of the following input with up to {max_words} words.
3 This summary should contain the information required to predict the students answer to any new question.
4 {cot_instruction}
5
6 Input:
7 {input_text}
8
9 Once again, please provide a **concise** summary of the students knowledge state.
10 {cot_instruction}
11 Keep your summary under {max_words} words.
Listing 1: Base Prompt for the Encoder.
1template: |
2 Here is some information about a student:
3 {bottleneck}
4
5 {y_prompt}
Listing 2: Base Prompt for the Decoder.
1template: |
2 Please produce a concise summary of the following input with up to {max_words} words.
3 This summary should contain the information required to predict the students answer to any new question.
4
5 Make sure to mention any misconception held by the student.
6 {cot_instruction}
7
8 Input:
9 {input_text}
10
11 Once again, please provide a **concise** summary of the students knowledge state.
12 {cot_instruction}
13 Keep your summary under {max_words} words.
Listing 3: Base Prompt for the Encoder when steering for mentioning "misconceptions" (Section B.4.1).
D.5 Code and hardware
Experiments were ran on four NVIDIA A100 GPUs with 80GB VRAM each, and 880GB of total RAM.